# Permutation Learning with Only N Parameters: From SoftSort to Self-Organizing Gaussians

### v1

@@ Abstract
Sorting and permutation learning are key concepts in optimization and machine learning, especially when organizing high-dimensional data into meaningful spatial layouts. The Gumbel-Sinkhorn method, while effective, requires parameters to determine a full permutation matrix, making it computationally expensive for large datasets. Low-rank matrix factorization approximations reduce memory requirements to (with ), but they still struggle with very large problems. SoftSort, by providing a continuous relaxation of the argsort operator, allows differentiable 1D sorting, but it faces challenges with multidimensional data and complex permutations. In this paper, we present a novel method for learning permutations using only parameters, which dramatically reduces storage costs. Our approach builds on SoftSort, but extends it by iteratively shuffling the indices of the elements to be sorted through a separable learning process. This modification significantly improves sorting quality, especially for multidimensional data and complex optimization criteria, and outperforms pure SoftSort. Our method offers improved memory efficiency and scalability compared to existing approaches, while maintaining high-quality permutation learning. Its dramatically reduced memory requirements make it particularly well-suited for large-scale optimization tasks, such as ”Self-Organizing Gaussians”, where efficient and scalable permutation learning is critical.
@@
While our proposed permutation learning method has broad applicability across various fields, this paper focuses on its use in distance-preserving grid layout algorithms for visual image sorting and 3D scene reconstruction in 3D Gaussian Splatting [7]. Sorting images based on similarity enhances visual perception, enabling users to efficiently explore hundreds of images simultaneously. In 3D scene reconstruction, organizing each scene attribute into a sorted 2D grid improves spatial correlation, and the resulting high-dimensional maps can be compressed using standard codecs, achieving substantial reductions without sacrificing rendering quality.
@@ I-B Heuristic Distance-Preserving Grid Layout Algorithms
Distance-Preserving Grid Layouts optimize the assignment of objects or images to grid positions by aligning spatial proximity with the similarity relationships of their feature vectors. Various algorithms have been proposed for arranging vectors on a 2D grid, most of which rely on heuristic methods that offer efficient performance. However, recent research [2] has demonstrated that gradient-based learning techniques can produce sorted grid layouts with superior sorting quality compared to traditional approaches. In the following sections, we present a brief overview of the most commonly used non-learnable methods.
@@
Dimensionality reduction techniques like t-SNE [19] or UMAP [10] can be applied to project high-dimensional input vectors onto the 2D plane before rearranging them on the grid. Fast placement algorithms for this task are discussed in [5]. Linear assignment solvers, such as the Jonker-Volgenant Algorithm [6], can map these positions to the best grid locations.
@@ I-C Contributions
ShuffleSoftSort requires only parameters. By using an iterated shuffling of the sorting indices, it iteratively reorders elements along different axes, enabling large-scale swaps not possible with SoftSort.
We show that ShuffleSoftSort achieves comparable performance to low-rank methods such as Kissing to Find a Match at reduced memory requirements.
@@ II Proposed Method
Figure 2 illustrates the principle of permutation learning with the goal of estimating a permutation matrix that sorts the input vectors into . Since is not differentiable, a continuous version is optimized by adjusting the weights based on a given loss function. The loss function ensures (a binarized version of ) forms a valid permutation while satisfying the optimization requirements.
@@
Our new approach is based on SoftSort, which was originally developed as a continuous relaxation for the argsort operator. It is defined as
with being the distance matrix between the sorted and the unsorted vector . Row-wise softmax increases matrix elements at positions where elements coincide, while suppressing the others. With decreasing the matrix converges to the permutation matrix.
@@
Figure 3 shows the problem using a 1D toy example. If the colors yellow and magenta were swapped, the color sorting according to minimum neighbor distance would be perfect. However, since reaching the target positions would require passing through the other colors, gradient-based learning would fail to find this path, as the quality would initially degrade. The same issue arises in Figure 1 when attempting to swap elements that are separated by multiple lines.
@@
Our solution is to iteratively reorganize/shuffle the indices of the data to be sorted and then apply SoftSort using the updated one-dimensional indices. Rearranging the 1D order allows elements to be repositioned that would not have been possible in the original order. For 2D grid-based sorting, a practical approach is to map indices into a two-dimensional structure and alternate sorting along horizontal and vertical axes, enabling more substantial adjustments than SoftSort alone. Figure 4 shows the principle of ShuffleSoftSort for the alternating 2D organization of the indices.
This strategy extends to other permutation learning tasks, with the key idea of continuously modifying the index order to allow for greater flexibility in sorting. However, simply rearranging the data and performing repeated sorting on the shuffled indexing does not work, because each new SoftSort iteration would cancel the effect of the previous one. To prevent this, the weights of the next sorting step must be initialized in a way that initially preserves the previous order. This can be achieved by setting the weights of the new SoftSort sorting in ascending linear order. The extent to which the old order is retained depends on the value of . During training, must be continuously decreased to successively reduce the influence of the neighboring elements. Algorithm 1 describes the ShuffleSoftSort sorting process.
For a memory-efficient implementation, it is crucial to compute the permutation matrix and the elements of the loss function, which would otherwise require elements, in a line-wise manner.
@@
The regularization parameters and were set to 1 and 2 respectively. SoftSort typically produces a valid permutation matrix. In rare cases, where the columns of the permutation matrix contain duplicates, is decreased and the SoftSort iterations are extended until a valid permutation is achieved.
@@ III Experimental Evaluation
Memory | Runtime [s] | Quality | |

### v2

@@ Abstract
Sorting and permutation learning are key concepts in optimization and machine learning, especially when organizing high-dimensional data into meaningful spatial layouts. The Gumbel-Sinkhorn method, while effective, requires parameters to determine a full permutation matrix, making it computationally expensive for large datasets. Low-rank matrix factorization approximations reduce memory requirements to (with ), but they still struggle with very large problems. SoftSort, by providing a continuous relaxation of the argsort operator, allows differentiable 1D sorting, but it faces challenges with multidimensional data and complex permutations. In this paper, we present a novel method for learning permutations using only parameters, which dramatically reduces storage costs. Our method extends SoftSort by iteratively shuffling the indices of the elements and applying a few SoftSort optimization steps per iteration. This modification significantly improves sorting quality, especially for multidimensional data and complex optimization criteria, and outperforms pure SoftSort. Our method offers improved memory efficiency and scalability compared to existing approaches, while maintaining high-quality permutation learning. Its dramatically reduced memory requirements make it particularly well-suited for large-scale optimization tasks, such as ”Self-Organizing Gaussians”, where efficient and scalable permutation learning is critical.
@@
While our proposed permutation learning method has broad applicability across various fields, this paper focuses on its use in distance-preserving grid layout algorithms for visual image sorting and 3D scene reconstruction in 3D Gaussian Splatting [7]. Sorting images based on similarity enhances visual perception, enabling users to efficiently explore hundreds of images simultaneously. In 3D scene reconstruction, organizing each scene attribute into a sorted 2D grid increases spatial correlation, and the resulting high-dimensional maps can be compressed using standard codecs, achieving substantial reductions without sacrificing rendering quality.
@@ I-B Heuristic Distance-Preserving Grid Layout Algorithms
Distance-Preserving Grid Layouts optimize the assignment of objects or images to grid positions by aligning spatial proximity with the similarity relationships of their feature vectors. Various algorithms have been proposed for arranging vectors on a 2D grid, most of which rely on heuristic methods that offer efficient performance. However, recent research [2] has demonstrated that gradient-based learning techniques can produce sorted grid layouts with superior sorting quality compared to traditional approaches. In the following sections, we present a brief overview of the most commonly used non-learning methods.
@@
Dimensionality reduction methods like t-SNE [19] and UMAP [10] can project high-dimensional vectors onto a 2D plane before grid arrangement. Linear assignment solvers, such as the Jonker–Volgenant algorithm [6], then map these positions to optimal grid locations. Fast placement strategies are discussed in [5].
@@ I-C Contributions
ShuffleSoftSort requires only parameters. By using an iterated shuffling of the sorting indices, it iteratively reorders elements along different pathes, enabling large-scale swaps not possible with SoftSort.
We show that ShuffleSoftSort achieves better performance to low-rank methods such as Kissing to Find a Match at reduced memory requirements.
@@ II Proposed Method
Figure 2 illustrates the principle of permutation learning with the goal of estimating a permutation matrix that sorts the input vectors into . Since is not differentiable, a continuous version is optimized by adjusting the weights based on a given loss function. The loss function enforces that , a binarized version of , constitutes a valid permutation while meeting the optimization requirements.
@@
Our new approach is based on SoftSort, which was originally developed as a continuous relaxation for the argsort operator, defined as
with being the L1 distance matrix between the sorted and the unsorted elements of the vector . Row-wise softmax increases matrix elements at positions where elements coincide, while suppressing the others. With decreasing the matrix converges to the permutation matrix.
@@
Figure 3 illustrates the SoftSort problem using a 1D toy example. If the yellow and magenta colors were swapped, the color ordering based on minimal neighbor distances would be perfect. However, reaching the correct positions would require passing through intermediate colors, causing gradient-based optimization to fail, as the quality initially degrades. A similar issue occurs in Figure 1, where swapping elements separated by multiple rows encounters the same problem.
@@
Our solution is to iteratively reorganize/shuffle the indices of the data to be sorted and then apply SoftSort using the updated one-dimensional indices. Rearranging the 1D order allows elements to be repositioned that would not have been possible in the original order.
Figure 4 illustrates the core concept of ShuffleSoftSort, highlighting how the alternating shuffled index organization improves sorting flexibility. Algorithm 1 describes the ShuffleSoftSort procedure for sorting the elements of a vector with shape .
The loss function is computed on the reverse-shuffled elements. Simply shuffling and repeatedly applying SoftSort is ineffective, as each iteration would undo the previous sorting. To address this, the weights in each SoftSort iteration must be set in a manner that initially preserves the previous order. This can be achieved by initializing the weights in a linear ascending order, combined with a small initial temperature parameter . During training, is gradually reduced to gradually reduce the influence of more distant neighbor elements. The final achievable sorting quality increases with the number of iterations .
This new sorting strategy extends to other permutation learning tasks, with the key idea of continuously modifying the index order to allow for greater flexibility in sorting. For a memory-efficient implementation, it is crucial to compute the permutation matrix and the loss elements in a row-wise manner, as the complete storage would require elements.
@@
The regularization parameters and were set to 1 and 2 respectively. SoftSort typically produces a valid permutation matrix. In very rare cases, where the columns of the permutation matrix contain duplicates, the SoftSort iterations are extended until a valid permutation is achieved.
@@ III Experimental Evaluation
Method | Memory | Runtime [s] | Quality |